Hello, My name is Lavalle.
I need help on a couple of problems that have been stressing me out. I hope you can help me better understand them.
1. How many 0.0125 grain doses can be made from 3/8 grains of a drug?

2. A formula for 1000 tablets contains 11.5 grams of an antihistamine. How many grams of the antihistamine should be used to prepare 475 tablets?

3. If a syringe contains 28 mg of drug in 3 ml how many milligrams of the drug would a patient receive if 1.2 ml is administered?

PLEASE HELP IF YOU CAN.

Answer by stanbon(48535)

1. Convert the 3/8 gr. to 0.375 gr. by dividing 3 by 8
Now divide 0.375 gr. by 0.0125 gr to get 30 doses.
2. Set up a propotion as follows:
x/475 = 11.5/1000
x = 475(11.5)/1000
x = 5.46 gr. of anti histamine to make 475 tablets.
3.Again, set up a proportion as follows:
x/ 1.2 ml = 28 mg/ 3 ml
x = 1.2(28)/3)
x = 11.2 mg
Cheers,
stan H.